You have been allocated a class A network address of 29.0.0.0. You need to create at least 20 networks and each network will support a maximum of 160 hosts. Would the following two subnet masks Work?
255.255.0.0 and or 255.255.255.0
1
Expert's answer
2020-05-16T13:46:47-0400

From the above,

Mask 255.255.0.0 has 8 bits for the sub net, and 16 bits for the host.

8 bits will accommodate"=2^8=256" sub nets

16 bits will accommodate"=2^{16}=64000" hosts

Mask 255.255.255.0 has 16 bits for the sub net, and 8 bits for the host.

So, total possible hosts"=2^8-2=254"

Hence, both the given will work.